In a right triangle,find the length of the side not given.
a=_____ b=1 c=√10

Answer by ewatrrr(24785)


Hi
applying the Pythagorean Theorem
a^2 + 1^2 = sqrt(10)^2
a^2 = 10 - 1
a ^2 = 9
a = 3 (Note: solution a = -3 cannot be used for a length)